On January 4, 2019, Columbus Company purchased new equipment for $693,000 that had a useful life of four years and a salvage value of $53,000.

Required:
Prepare a schedule showing the annual depreciation and end-of-year accumulated depreciation for the first three years of the asset’s life under the straight-line method, the sum-of-the-years’-digits method, and the double-declining-balance method.

Analyze:
If the double-declining balance method is used to compute depreciation, what would be the book value of the asset at the end of 2020?

Complete this question by entering your answers in the tabs below. Straight Line Sum of Year Digits Double Declining AnalyzeStraight Line Sum of Year Digits Double Declining Analyze Prepare a schedule showing the annual depreciation and end-of-yearStraight Line Sum of Year Digits Double Declining Analyze - --- Prepare a schedule showing the annual depreciation and end-ofStraight Line supporter de Straight Line Sum of Year Digits Double Declining s Anyse Analyze If the double-declining balance
Answer #1

1)

STRAIGHT-LINE METHOD
Year Acquisition Cost Salvage Value Useful Life Annual Depreciation Accumulated Depreciation
2019 $693000 53000 4 years $160000 $160000
2020 $693000 53000 4 years 160000 (160000+160000)= 320000
2021 $693000 53000 4 years 160000 (320000+160000)= 480000

Depreciation expense= (Acquisition cost-Salvage value)/Useful life

= $(693000-53000)/4= $160000

2) Sum of Years' digit= 4+3+2+1= 10

Cost less salvage= $693000-53000= $640000

SUM-OF-THE-YEARS'-DIGIT METHOD
Year Fraction Cost Less Salvage Annual Depreciation Accumulated Depreciation
2019 4/10 $640000 (640000*4/10)= $256000 $256000
2020 3/10 640000 (640000*3/10)= 192000 (256000+192000)= 448000
2021 2/10 640000 (640000*2/10)= 128000 (448000+128000)= 576000

3) Depreciation rate= 100*2/4= 50%

DOUBLE-DECLINING-BALANCE METHOD
Year Beginning Book Value Rate Annual Depreciation Accumulated Depreciation
2019 $693000 50% (693000*50%)= $346500 $346500
2020 (693000-346500)= 346500 50% (346500*50%)= 173250 (346500+173250)= 519750
2021 (693000-519750)= 173250 50% (173250*50%)= 86625 (519750+86625)= 606375

4) Book value at the end of 2019= Beginning Book Value at the starting of 2019-Accumulated Depreciation at the end of 2020

= $693000-519750

= $173250